OpenStudy (katecc379):

Bonnie has a container in the shape of a rectangular pyramid. The formula for the surface area of the enclosed space is S = lw + 0.5Ph. Solve for P.

OpenStudy (katecc379):

P = S – lw – 0.5h P = S + lw + 0.5h P = The quantity S minus l times w all divided by 0.5 times h P = S divided by the quantity l times w plus 0.5 times h

OpenStudy (katecc379):

those are the choices

OpenStudy (vuriffy):

Well, you must get P to one side.

OpenStudy (katecc379):

yes i just cannot seem to get it to one of their choices

OpenStudy (vuriffy):

I surely can inform you the first two choices are incorrect.

OpenStudy (katecc379):

yeah i got that far haha

OpenStudy (vuriffy):

Well look at this:

OpenStudy (vuriffy):

S = lw + 0.5Ph -lw Which becomes: S - lw = 0.5Ph

OpenStudy (katecc379):

oh so then just divide by .5

OpenStudy (katecc379):

but then the h

OpenStudy (katecc379):

so s-lw/.5h

OpenStudy (vuriffy):

The values 0.5 and h are both multiplied toward the quantity P, so yes indeed, you must divide.

OpenStudy (vuriffy):

Sorry my wifi suddenly dropped.

OpenStudy (vuriffy):

Leaving you with: P = (S - lw)/(0.5h)

OpenStudy (vuriffy):

Which is the third choice.